OpenStudy (anonymous):

How do you solve this... 7(3-X) + 2x = -9

OpenStudy (anonymous):

do you know how to distribute through parenthesis ?

OpenStudy (anonymous):

take the number right before the parenthesis and multiply it by everything in the parenthesis. 7(3 - x) + 2x = -9 7 * 3 - 7 * x + 2x = -9 21 - 7x + 2x = -9 are you understanding so far ?

OpenStudy (anonymous):

when you want to participate, just let me know

OpenStudy (anonymous):

@wallenbaughz do u need more help? :)

OpenStudy (anonymous):

he/she is unresponsive, so I do not know if he/she even understands what I have done so far